Assume an apartment building contains 120 units. 60% are one-bedrooms that rent for $1000/month and 40% are two-bedrooms that rent for $1,500 per month. Vacancy is 10% of total income. What is the effective gross income?
Answer #1

Given,

Total units = 120

One bedrooms = 60% of 120 = 72

Two-bedrooms = 40% of 120 = 48

Total monthly income = 72*1000+48*1500 = $144000

Vacancy is 10% of total income

So, effective gross income = $144000*0.9*12 = $1555200